OpenStudy (anonymous):

A triangle has side lengths that are all integers. The length of one side is four times as long as the second side, and the third side has length 20. What is the largest possible perimeter of the triangle?

OpenStudy (anonymous):

let this circle has sides \[a,b,c\]\[a=4x,b=x,c=20\]

OpenStudy (anonymous):

\[P(x)=20+x+4x=20+5x\]

OpenStudy (campbell_st):

there is a bit of triangle geometry that says the sum of the shorter sides of a triangle must be greater than the longer side so if 20 is the longer side 4x + x > 20 or 5x > 20 resulting in x > 5 if 4x is the longer side then allply the same logic x + 20 > 4x or -3x > -20 or x < 20/3 perhaps this helps... in finding the largest perimeter given you know the max and min values

OpenStudy (anonymous):

the first one results in x>4 x<20/3=6.67 so the largest x is when it is x=6(integer) so \[P(6)=20+5(6)=50\]

OpenStudy (anonymous):

yeah its correct

OpenStudy (anonymous):

i mean \[\huge 4\le x \le 6.67,x \in Z\implies x_{max}=6\]